OpenStudy (anonymous):

Determine whether the system of linear equations has one and only one solution, infinitely many solutions, or no solution. x − 3y = −5 4x + 3y = 10

OpenStudy (anonymous):

a) try to solve it if you get to one single solution x=something y=something_else then it is a single solution. b) if you left with x=something*y+something_else that is infinite solutions c) if you left with x=something and x=something_else - this is a contradiction thus - no solutions

OpenStudy (anonymous):

i put no solotions the answer was wrong

OpenStudy (anonymous):

with the first equation you solve for x subtracting it on both sides

OpenStudy (anonymous):

This has unique solution x-3y=-5 4x+3y=10. Multiplying first one by 4, 4x-12y=-20. Subtracting from second. 15y=30,y=2. x=1

OpenStudy (anonymous):

lets do (a) solve it (from line 1) x - 3y = -5 thus x=3y-5 (from line 2) 4x + 3y = 10 thus 4(3y-5)+3y=10 thus 12y-20+3y=10 thus 15y=30 thus y = 2 back to x=3y-5 means x=3*2-5 x=1

OpenStudy (anonymous):

giselle3214 you understand what i did?

OpenStudy (anonymous):

ohkay so with this 2x − y = 5 and 3x+ y = −6 (first equation ) x = y-5 ( second equation ) ??

OpenStudy (anonymous):

line two confused me bratner

OpenStudy (anonymous):

the idea is to start with one line as if it was the only one and solve for one of the unknowns. i solved for x. then go to the second line and substitute x with what you found from the first line . you should be left with an equation with one unknown. solve it and you are through.

OpenStudy (anonymous):

ohhhhhh you have to substitue for the second line i thought you solve for it again

OpenStudy (anonymous):

here is example of "no solution" situation: x+y=10 2x+2y=30 tell me if you can't calculate and make sure there is actually no solution.

OpenStudy (anonymous):

i get it know then

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

it equals out to the same thing

OpenStudy (anonymous):

i specifically "cooked" the equations to have no solution. please write your calculations so i can see what is wrong.

OpenStudy (anonymous):

with the first one x and y equals to one so that would make the second one equals itself

OpenStudy (anonymous):

no this is incorrect. concentrate on the first line and solve for x. tell me what you get.

OpenStudy (anonymous):

the first line is x+y = 10

OpenStudy (anonymous):

you minus x on both sides. y=x-10

OpenStudy (anonymous):

actually you should get y=10-x

OpenStudy (anonymous):

it matters the order?

OpenStudy (anonymous):

the sign infront of the x and 10 matters because => 15 = 20-5 but not 15 = 5-20

OpenStudy (anonymous):

see my point?

OpenStudy (anonymous):

ohkay so it does. i see your point

OpenStudy (anonymous):

no go to the second line and substitute y for 10-x ( we found that y = 10-x) the second line was 2x+2y=30 . you will be left with equation with one unknown (x) - solve it and tell me what you get.

OpenStudy (anonymous):

2x+2(10-x)=30

OpenStudy (anonymous):

good. now you can solve it for x.

OpenStudy (anonymous):

2x=20-2x=30

OpenStudy (anonymous):

x=10, x= -15

OpenStudy (anonymous):

not really. 2x+2(10-x)=30 means that 2x+20-2x=30

OpenStudy (anonymous):

i don't understand how the second "equals" sign got into one equation.

OpenStudy (anonymous):

now you have 2x+20-2x=30 and that leads us to 20=30 (2x-2x is 0) which is not true ("contradiction"). which means that the set of equations that i gave you is absurd and has no solution.

OpenStudy (anonymous):

im sooooo confsued

OpenStudy (anonymous):

2x − y = 5 3x + y = −6

OpenStudy (anonymous):

would that be one solution?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

start with the first line and solve for say... y because it is multiplied only by (minus sign). what you get?

OpenStudy (anonymous):

y=2x+5

OpenStudy (anonymous):

that is wrong. 2x-y=5 / -2x from both sides / you get -y=5-2x /multiply by -1 both sides you get y=-5+2x / change order you get y=+2x-5

OpenStudy (anonymous):

-5?

OpenStudy (anonymous):

yep . if you multiply 5 by -1 you get 5*(-1)=-5

OpenStudy (anonymous):

i see you have trouble with multiplication by -1. it goes like this -y = 5 - 2x /multiply by -1 both sides means (-1)*(-y) = -1*(5-2x) /open paranthesis y = -1*5-1*(-2x) y = -5+2x y = 2x - 5

OpenStudy (anonymous):

the whole thing is to get rid of that minus sign that we get stuck with in front of y

OpenStudy (anonymous):

now go to the second line and substitute that y with 2x-5

OpenStudy (anonymous):

3x+y+-6 3x+y(2x-5)

OpenStudy (anonymous):

its an equation it should have an equal sign. and by substituting i mean "write 2x-5 instead of y"

OpenStudy (anonymous):

the idea is that we "already know what y is! y is 2x-5" that is how we get rid of all y in the second line.

OpenStudy (anonymous):

try again?